OpenStudy (anonymous):

What expression is the simplified form of sqrt a^13???? HELP!

OpenStudy (dumbcow):

\[\sqrt{a^{13}} = \sqrt{a^{12}a} = \sqrt{a^{12}}\sqrt{a} = \sqrt{(a^{6})^{2}}\sqrt{a} = a^{6}\sqrt{a}\]

OpenStudy (anonymous):

Can you explain how to do that? Thanks!